OpenStudy (anonymous):

Tom baked 2 cakes using a circular cake pan of diameter 10 inches. He plans to make a layered cake with one cake on top of the other with frosting only in between the two cakes and on top of the second cake. Maria baked 3 cakes using a circular cake pan of diameter 6 inches. She plans to make a layered cake with one cake on top of the other and frosting only in between the cakes and on top of the third cake. The height of frosting in both their cakes is the same.

OpenStudy (anonymous):

Who requires more frosting?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@yulima

OpenStudy (anonymous):

@DemolisionWolf @historylover

OpenStudy (anonymous):

tom

OpenStudy (anonymous):

we want to find how tall each cake is?

OpenStudy (anonymous):

"Tom baked 2 cakes using a circular cake pan of diameter 10 inches. He plans to make a layered cake with one cake on top of the other with frosting only in between the two cakes and on top of the second cake." Tom: 2Cakes 10inch Diameter 2Cakes + 2Frosting=height "Maria baked 3 cakes using a circular cake pan of diameter 6 inches. She plans to make a layered cake with one cake on top of the other and frosting only in between the cakes and on top of the third cake. Maria: 3Cakes 6inch Diameter 3Cakes + 3Frosting = height

OpenStudy (anonymous):

No who requires more frosting

OpenStudy (anonymous):

Maria. She has to frost an area of 56.52 square inches, and Tom has to frost an area of 31.40 square inches. Tom. He has to frost an area of 157 square inches, and Maria has to frost an area of 84.78 square inches. Maria. She has to frost an area of 339.12 square inches, and Tom has to frost an area of 157 square inches. Tom. He has to frost an area of 62.8 square inches, and Maria has to frost an area of 56.52 square inches.

OpenStudy (anonymous):

Options

OpenStudy (anonymous):

awh ^_^ ok! so Tom frosted 2 cakes right? so the area of one cake of tom is Area = pi*r^2 recal that r = diameter/2 can you find the area for 1 of tom's cake for me?

OpenStudy (anonymous):

3.14 * 5^2 = 78.5

OpenStudy (anonymous):

correct ^_^ so 78.5 is the area of one cake for tom. since tom frosts 2 cakes we do 2*78.5=? this will be the amount of area tom will frost, what is it?

OpenStudy (anonymous):

175 is what he needs..Now do the same for Maria?

OpenStudy (anonymous):

correct ^_^ so for maria, we also need to find the area of one cake. so, Area = pi*r^2, remember that r=diameter/2 what's the area for one cake of maria's then?

OpenStudy (anonymous):

28.26

OpenStudy (anonymous):

now multiply by 3 cakes?

OpenStudy (anonymous):

yep! ^_^

OpenStudy (anonymous):

So it will be B?

OpenStudy (anonymous):

I think they wrote it wrong

OpenStudy (anonymous):

Tom. He has to frost an area of 157 square inches, and Maria has to frost an area of 84.78 square inches.

OpenStudy (anonymous):

Tom: 157 = 2*(78.5) = (2) * pi*5^2 Maria: 84.8 = 3*28.26

OpenStudy (anonymous):

(I think I verified Tom's area wrong the frist time, but 157 is what it's suppose to be, i just doubled checked ^_^)

OpenStudy (anonymous):

ok cool thanks.. want to help me with another? ur really good

OpenStudy (anonymous):

yea, just @ me when you have it posted ^_^

OpenStudy (anonymous):

ok1